andre must water his cactus plant every 9 days he must also feed it every 56 days he watered and fed his cactus in how many days will he water and feed his cactus on the same day again
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

andre must water his cactus plant every 9 days he must also feed it every 56 days he watered and fed his cactus in how many days will he water and feed his cactus on the same day again
-------
Find the least common multiple of 9 and 56
-----
Factored Form:: 9 = 3^2
Factored Form:: 56 = 8*7 = 2^3*7
------
LCM =product of each prime factor to its highest power= 2^3*3^7=8*2187= 17496 days
--------
Cheers,
Stan H.
-----------

Answer by ikleyn(52803)

.
The correct answer is   9*56 = 504 days.